OpenStudy (anonymous):

What is the solution of log 3x - 2125 = 3 ? Choose one answer. a. x = 1 over 3 b. x = 1 c. x = 7 over 3 d. x = 4

OpenStudy (mertsj):

Add 2125 to both sides and then put the equation into its exponential form and solve for x.

OpenStudy (anonymous):

Crap sorry its supposed to be 3x-2 then 125=3

OpenStudy (mertsj):

\[\log_{3x-2} 125=3\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

yes!

OpenStudy (mertsj):

\[(3x-2)^3=125\]

OpenStudy (mertsj):

\[(3x-2)^3=5^3\]

OpenStudy (mertsj):

Can you get it from there?

OpenStudy (anonymous):

what do I do know

OpenStudy (mertsj):

\[2^5=x^5\] What would x be?

OpenStudy (anonymous):

2?

OpenStudy (mertsj):

\[2^5=(2x)^5\]

OpenStudy (mertsj):

What would 2x be?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

I have no clue..

OpenStudy (mertsj):

Here is the concept I am trying to get across to you: If you have two expressions that are equal and the exponents are the same, the only way they could be equal is if the bases are also the same.

OpenStudy (mertsj):

\[2^3=5^3\] Could that possible be true?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

NO!!

OpenStudy (anonymous):

so would it be c

OpenStudy (mertsj):

So if \[2^3=x^3\] x for sure has to be what?

OpenStudy (mertsj):

Yes